OpenStudy (anonymous):

Use the quadratic formula to solve the equation. x^2-x=-10. What is the solution set?

OpenStudy (anonymous):

\[x^{2} - x + 10 = 0\] Plugging into the quadratic formula you get \[\frac{1 \pm \sqrt{39}i}{2}\]

OpenStudy (anonymous):

\[x=(1/2) \pm (i/2)\sqrt{39}\]